# UPCORE: Utility-Preserving Coreset Selection for Balanced Unlearning

### v1

@@ Abstract
User specifications or legal frameworks often require information to be removed from pretrained models, including large language models (LLMs). This requires deleting or “forgetting” a set of data points from an already-trained model, which typically degrades its performance on other data points. Thus, a balance must be struck between removing information and keeping the model’s other abilities intact, with a failure to balance this trade-off leading to poor deletion or an unusable model. To this end, we propose UPCORE (Utility-Preserving Coreset Selection), a method-agnostic data selection framework for mitigating collateral damage during unlearning. Finding that the model damage is correlated with the variance of the model’s representations on the forget set, we selectively prune the forget set to remove outliers, thereby minimizing model degradation after unlearning. We evaluate UPCORE across three standard unlearning methods consistently achieving a superior balance between the competing objectives of deletion efficacy and model preservation. To better evaluate this trade-off, we introduce a new metric, measuring the area-under-the-curve (AUC) across standard metrics. We find that UPCORE improves both standard metrics and AUC, benefitting from positive transfer between the coreset and pruned points while reducing negative transfer from the forget set to points outside of it. [^1] Our code and data is available at: https://github.com/Vaidehi99/UPCORE
Machine unlearning, dataset selection, LLM model editing
@@
1 Introduction
The widespread deployment of ML models, particularly large language models (LLMs), has raised significant concerns regarding data privacy, regulatory compliance, and ethical AI practices. These models are often trained on vast amounts of uncurated data scraped from the internet, inherently capturing sensitive, copyrighted, or undesirable content (Shokri et al., 2017; Carlini et al., 2019). As regulations like the European Union’s General Data Protection Regulation (GDPR) and the California Consumer Privacy Act (CCPA) empower individuals with the “right to be forgotten”, the need for efficient techniques that remove specific data or topics from trained models has become increasingly critical. Machine unlearning has emerged as a promising solution, enabling the targeted removal of data, concepts, or facts without the computational expense of retraining from scratch. Moreover, machine unlearning has benefits beyond compliance, addressing broader challenges such as mitigating harmful outputs, preserving intellectual property rights, and aligning LLMs with ethical and societal expectations (Jang et al., 2023). These practical uses have spurred growing interest in understanding, rethinking, and improving model editing and unlearning methodologies (Liu et al., 2024; Hase et al., 2024).
Given the growing adoption of LLMs, past work has proposed methods for developing and evaluating techniques for removing knowledge or skills from LLMs (Cao & Yang, 2015; Bourtoule et al., 2021; Nguyen et al., 2022) and steering their behavior in targeted ways (Sinitsin et al., 2020; Meng et al., 2022). However, these editing methods often involve modifying the model in ways that lead to unintended consequences, such as reduced model utility on tasks or knowledge unrelated to the target. Thus, a balance must be struck between the successful unlearning of undesired information and maintaining the utility of the model by minimizing collateral damage. For this reason, current approaches generally evaluate unlearning or model editing methods based on their efficacy in altering intended knowledge – measured by how successfully a “forget set” is removed – and based on their collateral effects on unrelated behaviors, i.e. the accuracy on a “retain set”. This kind of evaluation is especially crucial in realistic settings where unlearning happens on a topic. Such topic-based unlearning commonly arises in practical scenarios, for example, when deleting all information associated with an individual or with sensitive domains (Li et al., 2024). Here, the likelihood of over-generalization to similar topics beyond the domain being deleted is high.
A key gap in existing research lies in understanding the specific data characteristics that drive overgeneralization and collateral effects during unlearning. While prior work (Sheshadri et al., 2024; Chowdhury et al., 2024) has measured damage resulting from unlearning, it does not investigate how attributes of the data – such as its variance – contribute to collateral damage or whether these attributes can be controlled to optimize the trade-off between deletion efficacy and utility retention. Focusing on a topic-based setting where the forget set comprises semantically coherent groups of information, we seek to address these questions:
@@ 1 Introduction
We investigate which properties of the forget data correlate with collateral damage during unlearning. Our analysis reveals a strong positive correlation between the variance of the model’s hidden states corresponding to datapoints in the forget set (hidden state variance, or HSV), and the extent of collateral damage to the model after unlearning. In other words, unlearning a set of widely-distributed datapoints (as shown in Figure 1 (left)) leads to more damage than unlearning a more densely-distributed set. Building on this insight, we hypothesize that selectively curating a coreset with lower variance from the larger forget set can help optimize this trade-off, as shown in Figure 1 (right).
@@
To this end, we introduce UPCORE, which constructs a core forget set by systematically identifying and pruning data points in the forget set that contribute most to the variance and thereby to collateral damage. UPCORE organizes points into an Isolation Forest (Liu et al., 2008), which identifies anomalous points in a set. By pruning these points, we reduce the variance within the forget set, which we find leads to less damage. Crucially, in addition to reducing collateral damage, UPCORE in fact leverages it by identifying two separate kinds of collateral effects: (1) Negative collateral damage: Unintended degradation of unrelated model capabilities and (2) Positive collateral transfer: The intended impact on pruned data points removed to form the core forget set. This is illustrated in Figure 1, where pruned outlier points are still unlearned – despite not being a part of the coreset used for unlearning – due to positive transfer, and is further highlighted by our results in Table 1 and Table 3, which show that UPCORE results in better unlearning than a randomly selected subset while also having better knowledge retention on non-forget data. Moreover, UPCORE is method-agnostic: by focusing solely on the data, UPCORE can be applied to any data-driven unlearning framework.
@@
We evaluate UPCORE in prompt completion and question-answering settings and across three standard unlearning methods: Gradient Ascent (Jang et al., 2023), Refusal (Ouyang et al., 2022) and Negative Preference Optimization (NPO) (Zhang et al., 2024c), applying each unlearning algorithm directly to the optimized core forget set obtained using UPCORE, rather than the entire forget set. We measure three critical dimensions: (1) Unlearning effectiveness, measured by the successful removal of targeted knowledge in the forget set, paraphrased versions of removed information as well as prompts attempting to jailbreak the model.; (2) Unintended damage, where we quantify collateral effects on unrelated model capabilities; and (3) Intended transfer, where we analyze the impact on the pruned data points that were removed from the core forget set. While we follow past unlearning work in the metrics we use to measure the trade-off between the competing objectives, we also note that the current suite of metrics measures performance at a fixed point during unlearning. This can make comparisons across methods hard, as the trade-off between deletion efficacy and model utility varies across unlearning epochs. To address this, in addition to showing improvements on standard metrics, we introduce a novel set of metrics that report the area-under-the-curve (AUC) for the standard unlearning metric suite, reporting not just the performance at one fixed timestep, but measuring how a method trades off deletion with model utility across checkpoints (see Figure 3).
@@
Empirically, we find that across all three unlearning methods, UPCORE consistently has the highest AUC compared to baselines of unlearning on the complete forget set and choosing a random subset of forget points. In other words, UPCORE forms a Pareto frontier, doing a better job of maximizing unlearning effectiveness while also minimizing model damage. Moreover, UPCORE positively leverages generalization by transferring unlearning from the core set to the high-variance outlier points that were removed from the core forget set. Notably, it consistently outperforms baselines across all unlearning methods; this holds true when comparing AUC across multiple metrics (e.g. ROUGE on a “retain” set, on neighborhood data closely related to the forget set but not in it, etc.). UPCORE’s superior trade-off effectively generalizes to variations of the forgotten information, performing well on paraphrased versions of forgotten prompts as well as prompts intended to jailbreak the model. We also see these gains reflected in static evaluations of one checkpoint (as opposed to AUC, which evaluates across checkpoints); here, UPCORE obtains lower (better) ROUGE on the forget set than the random baseline while simultaneously incurring less model damage than the random and complete baselines, with the best (highest) ROUGE across all data not in the forget set.
@@ 2 Background and Related Work
Unlearning Methods for LLMs.
Machine unlearning is broadly categorized into exact unlearning, which ensures the model is indistinguishable from one retrained without the forget data, and approximate unlearning, which efficiently modifies existing model parameters to approximate this effect without full retraining. Due to the cost of retraining LLMs, most unlearning applied to LLMs (including ours) falls into the second category. One such approach trains the model to output an uninformative response instead of knowledge in the “forget set” via RLHF (Ouyang et al., 2022), maximizing the probability of a predefined response like, “I don’t know” (Wen et al., 2024). Yao et al. (2023) introduce a gradient ascent-based method for unlearning harmful content, outputting whitespace for harmful prompts. While effective in reducing harmful responses, this approach leads to notable performance degradation on normal prompts, underscoring the need for balance. Chen & Yang (2023) propose an unlearning framework that introduces an unlearning layer, demonstrating success in both classification and generation tasks with less extreme performance degradation. Eldan & Russinovich (2023) explore a novel network architecture designed to specifically unlearn copyrighted content in LLMs. On the evaluation side, Maini et al. (2024) present a benchmark for evaluating unlearning, which we use in our evaluation. Despite these advancements, balancing the trade-off between forgetting accuracy and preserving model utility remains an open challenge, motivating the need for more data-driven solutions. Our work addresses this gap by proposing a principled framework that focuses on minimizing collateral damage through data selection.
Model Editing for Unlearning.
Model editing provides an alternative approach to unlearning by directly modifying model weights to forget target facts (De Cao et al., 2021; Dai et al., 2022; Mitchell et al., 2022; Meng et al., 2022). This method aligns with privacy requirements (Zhang et al., 2024a), avoids data-side interventions (Debenedetti et al., 2024), and protects against white-box extraction attacks. Following model editing work like Patil et al. (2024b), our framework employs LoRA-based weight updates for controlled unlearning via standard unlearning objectives (See Section B.6 for more details).
Coreset Selection.
Coreset selection identifies representative subsets that preserve key dataset properties, improving computational efficiency. Given the NP-hard complexity of the exhaustive search, methods have focused on optimizing coverage, diversity, or importance (Sener & Savarese, 2018; Tan et al., 2023). By recognizing unequal contributions of data points, coreset selection has proven effective in supervised learning (Wei et al., 2015; Killamsetty et al., 2021b, a), enabling efficient performance. Our work forms new connections between these methods and the problem of unlearning in LLMs, where preserving utility and minimizing collateral damage are critical.
@@ 3.1 Problem Definition
To address this issue, we aim to construct a pruned forget set $D_{C}\subset D_{F}$, henceforth referred to as the core forget set, by removing points in $D_{F}$ that disproportionately drive over-generalization. The goal is to balance two competing objectives: (i) minimizing negative collateral damage, i.e., performance degradation on $D\setminus D_{F}$, and (ii) maintaining deletion accuracy, i.e., ensuring that the unlearning $D_{C}$ effectively deletes the undesirable knowledge associated with the original forget set $D_{F}$. More formally, given a damage metric $\text{Damage}_{(\mathcal{U},D_{C})}(M,M^{\prime},D\setminus D_{F})$ that quantifies the impact of unlearning $D_{C}$ on $D\setminus D_{F}$, and a deletion accuracy metric $\text{DelAcc}_{(\mathcal{U},D_{C})}(M^{\prime},D_{F})$ that evaluates the effectiveness of $\mathcal{U}$ in forgetting $D_{F}$ after unlearning on $D_{C}$, the problem can be formulated as an optimization task:
$D_{C}=\arg\min_{D_{C}\subseteq D_{F}}$ $\bigg{(}\text{Damage}_{(\mathcal{U},D_{C})}\big{(}M,M^{\prime},(D\setminus D_{F})\big{)}$
$-\lambda\cdot\text{DelAcc}_{(\mathcal{U},D_{C})}\big{(}M^{\prime},D_{F}\big{)}\bigg{)}$
where $\lambda>0$ is a hyperparameter that controls the trade-off between the competing objectives.
@@ 3.2 Variance as a Measure of Collateral Damage
Building on prior work analyzing the cross-task generalization of forgetting methods (Zhang et al., 2024b), we investigate the relationship between attributes of the forget set $D_{F}$ and their impact on collateral damage during unlearning i.e. $\text{Damage}_{(\mathcal{U},D_{C})}\big{(}M,M^{\prime},(D\setminus D_{F})\big{)}$. Specifically, we identify the variance $Var(D_{F})$ as a critical predictor of overgeneralization. To systematically evaluate this relationship, we analyze question-answer (QA) pairs generated from Wikipedia documents across diverse topics. Each topic-specific dataset acts as the forget set $D_{F}$, with unlearning applied sequentially, one topic at a time. For each forget set, we compute variance using the hidden states of the last token and the penultimate layer of the question. We then compute retain set performance as the model utility metric proposed by Maini et al. (2024), which measures performance on preserved data points after unlearning. The results, visualized in Figure 7(a) in Section B.2, demonstrate a strong negative correlation between HSV and model utility, indicating that variance is a potential driver of over-generalization. These findings underscore the importance of identifying and excluding points that lead to higher variance i.e. outliers to mitigate utility loss. In Section B.2 we show similar analyses for other attributes such as model confidence and gradient similarity but find no strong correlation between utility degradation and these attributes.
@@ 3.3 UPCORE: Core Forget Set Selection
To achieve variance minimization in the forget set $D_{F}$, UPCORE frames the problem as an outlier detection task. UPCORE provides two key benefits: (1) It mitigates negative collateral damage by pruning outliers to form a more compact core forget set, and (2) It strategically exploits collateral over-generalization to extend unlearning beyond the core forget set, effectively removing the pruned points as well. As shown in Figure 1, what might traditionally be viewed as detrimental collateral damage – when it affects points outside the forget set ($D_{F}$) – can be turned to our advantage when it impacts untrained data points within the forget set that were pruned ($D_{F}\setminus D_{C}$).
@@
To detect these outliers, we use the Isolation Forest algorithm (Liu et al., 2008), an unsupervised learning technique that efficiently identifies anomalous data points. Isolation Forest works by recursively partitioning the dataset using random feature selections and random split values. Points that are isolated with fewer partitions, i.e. are isolated more easily, are considered outliers, as they differ significantly from the majority of the data. This makes the Isolation Forest algorithm particularly effective for high-dimensional data where traditional distance-based outlier detection methods may fail. These outliers, characterized by their isolation in the feature space, are likely to contribute to high variance and over-generalization during the unlearning process. UPCORE proceeds as follows (illustrated in Figure 2):
@@
Stage 1: Hidden Feature Extraction: We extract hidden state representations $\mathcal{H}$ from the model’s penultimate layer (See Figure 2 left), corresponding to the final token of each question in $D_{F}$. These representations, which reflect the model’s internal representation of the data, serve as input features for outlier detection. This step is guided by our analysis in Section 3.2, which highlights the strong link between hidden state variance and collateral damage.
@@
Stage 2: Training the Isolation Forest and Computing Anomaly Scores: We train an Isolation Forest model $\mathcal{I}$ on the forget set $D_{F}$ to model its distribution, recursively partitioning the data to detect outliers (see Figure 2 middle). Points isolated more quickly and requiring fewer splits are flagged as outliers, indicating disproportionate contributions to variance in the hidden state space. For each $d\in D_{F}$, $\mathcal{I}$ assigns an anomaly score $\text{score}(d)$ based on the average path length $h(d)$ required to isolate the point across an ensemble of binary trees. Shorter path lengths correspond to higher anomaly scores, indicating points that contribute to variance and thus collateral effects. Additional details are provided in Appendix A.
@@
Stage 3: Prune Outliers and Setting Stopping Criterion: To construct the pruned coreset $D_{C}$, we apply a threshold $\tau$ on the anomaly scores from the Isolation Forest model: $D_{C}=\{d\in D_{F}\mid\text{score}(d)\leq\tau\}$. Data points with scores above $\tau$ are excluded as outliers, as they disproportionately contribute to variance. We hypothesize that removing these outliers will reduce utility degradation while preserving core information for forgetting. The threshold $\tau$ is determined via a stopping criterion, which can be chosen as follows: (1) Coreset Size Control: Specify a desired coreset size $|D_{C}|$ to ensure an appropriate number of inliers (2) Proportional Pruning: Select the top $k\%$ of points with the lowest anomaly scores to maintain a consistent pruning ratio. By selecting $\tau$ based on user requirements, UPCORE provides fine-grained control over the trade-off, ensuring the construction of a robust coreset. In practice, we prune 10% of the data points in our main experiments and additionally conduct scaling experiments that vary the pruning percentage to analyze its impact on the trade-off dynamics.
@@
Stage 4: Unlearning on the Coreset: After selecting the pruned coreset $D_{C}$, UPCORE applies the unlearning algorithm $\mathcal{U}$ to the model $M$, resulting in $M^{\prime}_{UPCORE}=\mathcal{U}(M,D_{C})$ (See Figure 2 end). This process removes the influence of $D_{F}$ while minimizing utility degradation on $D\setminus D_{F}$. By focusing on $D_{C}$, our approach ensures targeted unlearning and positively leverages collateral effects, as unlearning $D_{C}$ also deletes much of $D_{F}$’s influence, even those parts not explicitly included in $D_{C}$.

### v2

@@ Abstract
User specifications or legal frameworks often require information to be removed from pretrained models, including large language models (LLMs). This requires deleting or “forgetting” a set of data points from an already-trained model, which typically degrades its performance on other data points. Thus, a balance must be struck between removing information and keeping the model’s other abilities intact, with a failure to balance this trade-off leading to poor deletion or an unusable model. To this end, we propose UPCORE (Utility-Preserving Coreset Selection), a method-agnostic data selection framework for mitigating collateral damage during unlearning. Finding that the model damage is correlated with the variance of the model’s representations on the forget set, we selectively prune the forget set to remove outliers, thereby minimizing model degradation after unlearning. Across three standard unlearning methods, UPCORE consistently achieves a superior balance between the competing objectives of deletion efficacy and model preservation. To better evaluate this trade-off, we introduce a new metric, measuring the area-under-the-curve (AUC) across standard metrics. Our results show that UPCORE improves both standard metrics and AUC, benefiting from positive transfer between the coreset and pruned points while reducing negative transfer from the forget set to points outside of it. [^1] Correspondence to: Vaidehi Patil ¡vaidehi@cs.unc.edu¿. Code and data: https://github.com/Vaidehi99/UPCORE
1 Introduction
The widespread deployment of ML models, especially large language models (LLMs), has raised concerns around privacy, regulation, and ethical use. Trained on massive, uncurated web data, these models often memorize sensitive, copyrighted, or undesirable content (Shokri et al., 2017; Carlini et al., 2019). With regulations like the GDPR and CCPA granting individuals the “right to be forgotten,” efficient methods for removing specific data or topics from pre-trained models are increasingly necessary. Machine unlearning offers a promising solution by enabling targeted data removal without full retraining. Beyond regulatory compliance, it also helps reduce harmful outputs, protect intellectual property, and align LLMs with societal values (Jang et al., 2023). These challenges have driven renewed interest in improving model editing and unlearning techniques (Liu et al., 2024; Hase et al., 2024).
@@ 1 Introduction
Given the growing use of LLMs, prior work has proposed methods for removing knowledge or skills (Cao and Yang, 2015; Bourtoule et al., 2021; Nguyen et al., 2022) and steering behavior in targeted ways (Sinitsin et al., 2020; Meng et al., 2022). However, such editing often induces unintended side effects, reducing utility on unrelated tasks. Effective unlearning therefore requires balancing deletion of undesired information with preservation of overall model utility. To evaluate this, current methods assess both deletion success (via “forget set” performance) and collateral effects on unrelated behaviors (via “retain set” accuracy). This is particularly important in realistic, topic-level unlearning scenarios—such as removing information about an individual or sensitive domain (Li et al., 2024)—where deletion may cause over-generalization and degrade performance on semantically or structurally similar inputs.
A key gap in existing research lies in understanding the specific data characteristics that drive over-generalization and collateral effects during unlearning. While prior work (Sheshadri et al., 2024; Chowdhury et al., 2024) has measured damage resulting from unlearning, it does not investigate how attributes of the data – such as its variance – contribute to collateral damage or whether these attributes can be controlled to optimize the trade-off between deletion efficacy and utility retention. Focusing on a topic-based setting where the forget set comprises semantically coherent groups of information, we seek to address these questions:
@@
We investigate which properties of the forget data correlate with collateral damage during unlearning. Our analysis reveals a strong positive correlation between the variance of the model’s hidden states corresponding to datapoints in the forget set (hidden state variance, or HSV), and the extent of collateral damage to the model after unlearning. In other words, unlearning a set of widely-distributed datapoints (as shown in Fig. 1 (left)) leads to more damage than unlearning a more densely-distributed set. Building on this insight, we hypothesize that selectively curating a coreset with lower variance from the larger forget set can help optimize this trade-off, as shown in Fig. 1 (right).
@@
To this end, we introduce UPCORE, which constructs a core forget set by systematically identifying and pruning data points in the forget set that contribute most to the variance and thereby to collateral damage. UPCORE organizes points into an Isolation Forest (Liu et al., 2008), which identifies anomalous points in a set. By pruning these points, we reduce the variance within the forget set, which we find leads to less damage. Crucially, in addition to reducing collateral damage, UPCORE in fact leverages it by identifying two separate kinds of collateral effects: (1) Negative collateral damage: Unintended degradation of unrelated model capabilities and (2) Positive collateral transfer: The intended impact on pruned data points removed to form the core forget set. This is illustrated in Fig. 1, where pruned outlier points are still unlearned – despite not being a part of the coreset used for unlearning – due to positive transfer, and is further highlighted by our results in Table 1 and Table 3, which show that UPCORE results in better unlearning than a randomly-selected subset while also having better knowledge retention on non-forget data. We show positive collateral transfer enabled by UPCORE in Table 4, with deletion transferring to points outside the coreset. Moreover, our focus on data makes UPCORE method-agnostic: it can be applied to any data-driven unlearning framework.
@@
We evaluate UPCORE in prompt completion and question-answering settings and across three standard unlearning methods: Gradient Ascent (Jang et al., 2023), Refusal (Ouyang et al., 2022; Maini et al., 2024) and Negative Preference Optimization (NPO) (Zhang et al., 2024b), applying each unlearning algorithm directly to the optimized core forget set obtained using UPCORE, rather than the entire forget set. We measure three critical dimensions: (1) Unlearning effectiveness, measured by the successful removal of targeted knowledge in the (a) forget set, (b) paraphrased versions of removed information as well as (c) prompts attempting to jailbreak the model.; (2) Unintended damage, where we quantify collateral effects on unrelated model capabilities; and (3) Intended transfer, where we analyze the impact on the pruned data points that were removed from the core forget set as shown Section 5.2. While we follow past unlearning work in the metrics we use to measure the trade-off between the competing objectives, we also note that the current suite of metrics measures performance at a fixed point during unlearning. This can make comparisons across methods hard, as the trade-off between deletion efficacy and model utility varies across unlearning gradient update steps. To address this, in addition to showing improvements on standard metrics, we introduce a novel set of metrics that report the area-under-the-curve (AUC) for the standard unlearning metric suite, reporting not just the performance at one fixed timestep, but measuring how a method trades off deletion with model utility across checkpoints (see Fig. 4).
@@
Empirically, we find that across all three unlearning methods, UPCORE consistently has the highest AUC compared to baselines of unlearning on the complete forget set and choosing a random subset of forget points. In other words, UPCORE forms a Pareto frontier, maximizing unlearning effectiveness while also minimizing model damage. Moreover, UPCORE positively leverages generalization by transferring unlearning from the core set to the high-variance outlier points that were removed from the core forget set. Notably, it consistently beats baselines across all unlearning methods; this holds true across multiple metrics (e.g. ROUGE on a “retain” set, on neighborhood data closely related to (but not in) the forget set, etc.). UPCORE’s superior trade-off effectively generalizes to variations of the forgotten information, performing well on paraphrased versions of forgotten prompts as well as prompts intended to jailbreak the model. We also see these gains reflected in static evaluations of one checkpoint (as opposed to AUC, which evaluates across checkpoints); here, UPCORE obtains lower (better) ROUGE on the forget set than the random baseline while simultaneously incurring less model damage than the random and complete baselines, with the best (highest) ROUGE across all data not in the forget set.
@@ 2 Background and Related Work
Unlearning Methods for LLMs.   Machine unlearning methods fall into two categories: exact unlearning, which ensures the model is indistinguishable from one retrained without the forget data, and approximate unlearning, which modifies model parameters efficiently without full retraining. Due to the high cost of retraining LLMs, most approaches—including ours—use the latter. One class of methods trains models via RLHF to produce uninformative responses (e.g., “I don’t know”) on forget prompts (Ouyang et al., 2022; Wen et al., 2024). Yao et al. (2023) apply gradient ascent to suppress harmful outputs, substituting them with whitespace, though this causes utility loss on benign prompts. To mitigate such degradation, Chen and Yang (2023) introduce an unlearning layer effective across tasks, while Eldan and Russinovich (2023) propose a specialized architecture for removing copyrighted content. For evaluation, Maini et al. (2024) present a benchmark we adopt. Despite these advances, managing the trade-off between unlearning and utility remains challenging. We address this by introducing a data-driven coreset selection framework to minimize collateral damage.
Model Editing for Unlearning.   Model editing provides an alternative approach to unlearning by directly modifying model weights to forget target facts (De Cao et al., 2021; Dai et al., 2022; Mitchell et al., 2022; Meng et al., 2022). Following model editing work (Patil et al., 2024b), our framework employs LoRA updates with standard unlearning objectives (See Section 8.8 for more details).
Coreset Selection.   Coreset selection identifies representative subsets that preserve key dataset properties, improving computational efficiency. Given the NP-hard complexity of the exhaustive search, methods have focused on optimizing coverage, diversity, or importance (Sener and Savarese, 2018; Tan et al., 2023). By recognizing unequal contributions of data points, coreset selection has proven effective in supervised learning (Wei et al., 2015; Killamsetty et al., 2021a, b), enabling efficient performance. Our work forms new connections between these methods and the problem of unlearning in LLMs, where preserving utility and minimizing collateral damage are critical.
@@ 3.1 Problem Definition
To address this issue, we aim to construct a pruned forget set $D_{C}\subset D_{F}$, henceforth referred to as the core forget set, by removing points in $D_{F}$ that disproportionately drive over-generalization. The goal is to balance two competing objectives: (i) minimizing negative collateral damage, i.e., performance degradation on $D\setminus D_{F}$, and (ii) maintaining deletion accuracy, i.e., ensuring that the unlearning $D_{C}$ effectively deletes the undesirable knowledge associated with the original forget set $D_{F}$. More formally, given a damage metric $\text{Damage}_{(\mathcal{U},D_{C})}(M,M^{\prime},D\setminus D_{F})$ that quantifies the impact of unlearning $D_{C}$ on $D\setminus D_{F}$, and a deletion accuracy metric $\text{DelAcc}_{(\mathcal{U},D_{C})}(M^{\prime},D_{F})$ that evaluates the effectiveness of $\mathcal{U}$ in forgetting $D_{F}$ after unlearning on $D_{C}$, the problem can be formulated as an optimization task: $D_{C}=\arg\min_{D_{C}\subseteq D_{F}}\bigg{(}\text{Damage}_{(\mathcal{U},D_{C})}\big{(}M,M^{\prime},(D\setminus D_{F})\big{)}-\lambda\cdot\text{DelAcc}_{(\mathcal{U},D_{C})}\big{(}M^{\prime},D_{F}\big{)}\bigg{)}$ where $\lambda>0$ is a hyperparameter that controls the trade-off between the competing objectives.
@@ 3.2 Variance as a Measure of Collateral Damage
Building on prior work analyzing the cross-task generalization of forgetting methods (Zhang et al., 2024a), we investigate the relationship between attributes of the forget set $D_{F}$ and their impact on collateral damage during unlearning i.e. $\text{Damage}_{(\mathcal{U},D_{C})}\big{(}M,M^{\prime},(D\setminus D_{F})\big{)}$. Specifically, we identify the variance $Var(D_{F})$ as a critical predictor of overgeneralization. Since directly selecting optimal subsets based on unlearning performance is computationally infeasible – due to the high cost of repeatedly retraining or unlearning – our goal is to develop scalable heuristics like variance that approximate which subsets minimize collateral damage, enabling efficient and practical coreset selection for unlearning. To systematically evaluate the relationship between variance and unlearning, we analyze question-answer (QA) pairs generated from Wikipedia documents across diverse topics. Each topic-specific dataset acts as the forget set $D_{F}$, with unlearning applied separately, one topic at a time. For each forget set, we compute variance using the hidden states of the last token and the penultimate layer of the model. We then compute retain set performance as the model utility metric proposed by Maini et al. (2024), which measures performance on preserved data points after unlearning. The results, visualized in Fig. 8(a) in Section 5.4, demonstrate a strong negative correlation between HSV and model utility, indicating that variance is a potential driver of over-generalization. These findings underscore the importance of identifying and excluding points that lead to higher variance i.e. outliers to mitigate utility loss. In Section 5.4 we show similar analyses for other attributes such as model confidence and gradient similarity but find no strong correlation between utility degradation and these attributes.
@@ 3.3 UPCORE: Core Forget Set Selection
To achieve variance minimization in the forget set $D_{F}$, UPCORE frames the problem as an outlier detection task. UPCORE provides two key benefits: (1) It mitigates negative collateral damage by pruning outliers to form a more compact core forget set, and (2) It strategically exploits collateral over-generalization to extend unlearning beyond the core forget set, effectively removing the pruned points as well. As shown in Fig. 1, what might traditionally be viewed as detrimental collateral damage – when it affects points outside the forget set ($D_{F}$) – can be turned to our advantage when it impacts untrained data points within the forget set that were pruned ($D_{F}\setminus D_{C}$).
@@
To detect these outliers, we use the Isolation Forest algorithm (Liu et al., 2008), an unsupervised learning technique that efficiently identifies anomalous data points. Isolation Forest works by recursively partitioning the dataset using random feature selections and random split values. Points that are isolated with fewer partitions, i.e. are isolated more easily, are considered outliers, as they differ significantly from the majority of the data. This makes the Isolation Forest algorithm particularly effective for high-dimensional data where traditional distance-based outlier detection methods may fail. These outliers, characterized by their isolation in the feature space, are likely to contribute to high variance and over-generalization during the unlearning process. UPCORE proceeds as follows (illustrated in Fig. 2):
@@
Stage 1: Hidden Feature Extraction: We extract hidden state representations $\mathcal{H}$ from the model’s penultimate layer, as it typically encodes high-level semantic abstractions while retaining generalization capacity, unlike the final layer which is often biased toward task-specific outputs (Skean et al., 2025) (See Fig. 2 left), corresponding to the final token of each question in $D_{F}$. These representations, which reflect the model’s internal representation of the data, serve as input features for outlier detection. This step is guided by our analysis in Section 3.2, which highlights the strong link between hidden state variance and collateral damage.
@@
Stage 2: Training the Isolation Forest and Computing Anomaly Scores: We train an Isolation Forest model $\mathcal{I}$ on the forget set $D_{F}$ to model its distribution, recursively partitioning the data to detect outliers (see Fig. 2 middle). Points isolated more quickly and requiring fewer splits are flagged as outliers, indicating disproportionate contributions to variance in the hidden state space. For each $d\in D_{F}$, $\mathcal{I}$ assigns an anomaly score $\text{score}(d)$ based on the average path length $h(d)$ required to isolate the point across an ensemble of binary trees. Shorter path lengths correspond to higher anomaly scores, indicating points that contribute to variance and thus collateral effects. Additional details are provided in Section 7.
@@
Stage 3: Prune Outliers and Setting Stopping Criterion: To construct the pruned coreset $D_{C}$, we apply a threshold $\tau$ on the anomaly scores from the Isolation Forest model: $D_{C}=\{d\in D_{F}\mid\text{score}(d)\leq\tau\}$. Data points with scores above $\tau$ are excluded as outliers, as they disproportionately contribute to variance. We hypothesize that removing these outliers will reduce utility degradation while preserving core information for forgetting. The threshold $\tau$ is determined via a stopping criterion, which can be chosen as follows: (1) Coreset Size Control: Specify a desired coreset size $|D_{C}|$ to ensure an appropriate number of inliers (2) Proportional Pruning: Select the top $k\%$ of points with the lowest anomaly scores to maintain a consistent pruning ratio. By selecting $\tau$ based on user requirements, UPCORE provides fine-grained control over the trade-off. In practice, we prune 10% of the data points in our main experiments and additionally conduct scaling experiments that vary the pruning percentage (see Section 8.1) to analyze its impact on the trade-off dynamics.
@@
Stage 4: Unlearning on the Coreset: After selecting the pruned coreset $D_{C}$, UPCORE applies the unlearning algorithm $\mathcal{U}$ to the model $M$, resulting in $M^{\prime}_{\textsc{UPCORE}{}}=\mathcal{U}(M,D_{C})$ (See Fig. 2 end). This process removes the influence of $D_{F}$ while minimizing utility degradation on $D\setminus D_{F}$. By focusing on $D_{C}$, UPCORE ensures targeted unlearning and positively leverages collateral effects, as unlearning $D_{C}$ also deletes much of $D_{F}$’s influence, even those parts not explicitly included in $D_{C}$. We later show this positive transfer empirically in Table 4.